An educational psychologist wishes to know the mean number of words a third grader can read per minute. She wants to make an estimate at the 98% level of confidence. For a sample of 1116 third graders, the mean words per minute read was 28.1. Assume a population standard deviation of 4.2. Construct the confidence interval for the mean number of words a third grader can read per minute. Round your answers to one decimal place.
Answer #1

Solution :

Given that,

\bar x = 28.1

\sigma = 4.2

n = 1116

At 98% confidence level the z is ,

\alpha = 1 - 98% = 1 - 0.98 = 0.02

\alpha / 2 = 0.02 / 2 = 0.01

Z\alpha/2 = Z0.01 = 2.326

Margin of error = E = Z\alpha/2* (\sigma /\sqrtn) = 2.326 * (4.2 / \sqrt1116) = 0.3

At 98% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

28.1 - 0.3 < \mu < 28.1 + 0.3

27.8 < \mu < 28.4

(27.8 , 28.4)